Are 12 items that cost a total of $1.68,a better buy than 16 items that cost a total of $2.18?
Answer by jim_thompson5910(35256)

$1.68/12 = $0.14/1

So if 12 items cost $1.68, then each item costs $0.14 or 14 cents.

-------------------------------------------------------

$2.18/16 = 0.13625/1 = $0.14 (round up to the nearest penny)

So if 16 items cost $2.18, then each item costs $0.14 or 14 cents.

So the two deals are identical.